OpenStudy (anonymous):

What are the steps for this equation? 4x – 4 + 3x = 3x + 5 – 5x

OpenStudy (anonymous):

4x – 4 + 3x = 3x + 5 – 5x 4x+3x-4=3x-5x+5 ( put your common factors together) 7x-4=-2x +5 ( do the operations) 9x=9 x=1

OpenStudy (anonymous):

4x – 4 + 3x = 3x + 5 – 5x 4x+3x-4=3x-5x+5 ( put your common factors together) 7x-4=-2x +5 ( do the operations) 9x=9 ( we got this by adding four on both sides and adding 2x on both sides) x=19 divide 9 by both sides)